Find the area of the right triangle.
The sides are..height=6inches
base=7 inches
hypotenuse= 9 inches

THANK YOU

Found 2 solutions by solver91311, rothauserc:
Answer by solver91311(24713)



Your question makes no sense. 36 plus 49 = 85, but , so this is NOT a right triangle. If it is not a right triangle then none of the sides represent the height. So your statement: "The sides are. . height 6 inches" is the part that doesn't make sense.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by rothauserc(4718)

area (A) of triangle = 1/2 * base * height
A = 1/2 * 7 * 6 = 21 square inches